# Traffic Engineering in Large-scale Networks with Generalizable Graph Neural Networks

### v1

@@ V-B2 TEAL [39]
We used the open-source implementation of TEAL which solves TE with the same objective as TELGEN—maximizing the total flow. For fair comparison, we translated datasets generated for TELGEN into the same format as TEAL’s input, and compared their performance on the same datasets. As a remark, TEAL is designed in a way such that testing must be on the same topology and SD pairs as training.
@@ V-B3 HARP [3]
We used the open-source implementation of HARP, and translated our datasets into the input format of HARP. Note that HARP’s model architecture is specifically designed to minimize the MLU, and cannot generalize to other objectives such as the total flow maximization as in TEAL or TELGEN. To make sure the solutions are comparable under the same objective, we scale HARP’s traffic split ratios such that they fully respect the link capacities (i.e., MLU = $1$).
One issue of HARP is that it has a large memory footprint on large topologies, and even the latest hardware platform available to us (as below) could not execute HARP on the largest topology—the full ASN. To make a fair comparison, we trained and tested all three algorithms above on B4, and randomly sampled smaller subnetworks of ASN.
@@ V-B4 SciPy [8]
Hardware platform: All the algorithms were implemented and evaluated on a desktop server with Intel Xeon Gold 5317 CPU (3.0GHz, 48 cores), 256GB RAM and Nvidia A100 GPU (40GB). All experiments were run with $3$ random seeds and averages are shown below.
@@ V-C Evaluation Metrics
It is similar to the constraint loss but only considers final output of the algorithm. The objective gap without constraint violation (OnoCGap) is the objective gap when the solution is scaled by the maximum ratio of constraint violation, thereby ensuring the final solution is feasible. Finally, the training time per epoch and the prediction time measure the time spent on training for each epoch on average, and the time for making each prediction on average during testing (or running IPM on one instance till optimality for SciPy), respectively.
TABLE II: OnoCGap (%) for HARP, TEAL and TELGEN
@@
$\text{ASN}_{510}$ 3902.69 99.98 1.95
V-D Evaluation Results
V-D1 Comparing TELGEN, TEAL and HARP
We first compared the performance of TELGEN to TEAL and HARP, the state-of-the-art learning-based TE solutions. Since TEAL needs to be trained and tested on the same network topology, while HARP cannot execute on the whole ASN graph, we evaluated all three algorithms by training and testing them on datasets generated on B4 and two connected sub-networks of ASN with $98$ nodes and $202$ edges, $510$ nodes and $1600$ edges denoted as $\text{ASN}_{98}$ and $\text{ASN}_{510}$, respectively.
Table II shows OnoCGap—objective gap when each solution was rounded to be feasible—of TELGEN, TEAL and HARP on the three datasets. First, we observe that TELGEN achieved negligible ($<3\%$) objective gap, and the performance was consistent across topologies of different sizes. TELGEN strongly outperformed both TEAL and HARP, both of which had over $26\%$ gap on B4, and over $73\%$ gap on ASN datasets. We independently investigated the inferior performance of TEAL and HARP on our test sets. For TEAL, it was mainly because of the difference in traffic demands: TEAL’s original paper considered full traffic matrices with pairwise traffic demands, while our datasets included traffic demands only for randomly selected SD pairs. This led to TEAL inefficiently optimizing for many unnecessary SD pairs, showing its model’s inflexibility when facing diverse TE requirements. For HARP, the main reason was that it was designed for a different objective (MLU) than both TEAL and TELGEN, and could not be trivially modified to adopt a different objective due to its MLU-oriented model architecture. As such, its performance suffered when evaluated under the total flow objective since it unnecessarily constrained flows along non-bottlenecked paths.
@@ V-D1 Comparing TELGEN, TEAL and HARP
$\text{ASN}_{510}$ 20474.78 358.10 1.53
In Table III, both TEAL and HARP showed comparable or faster training time per epoch as TELGEN on B4, but both became at least 3×3$\times$3 × and 48×48$\times$48 × slower than TELGEN when facing the larger ASN98 and ASN510 topologies. For prediction time when making a prediction, TELGEN again significantly outperformed TEAL by over 3×3$\times$3 × for B4 and ASN98, 358×358$\times$358 × for ASN510; HARP by over 18×18$\times$18 × for B4, 500×500$\times$500 × for ASN98 and 13000×13000$\times$13000 × for ASN510 due to its simple, fully GPU-compatible architecture.
Especially, we observe that TELGEN kept almost a constant prediction time when scaling to drastically different network sizes. TELGEN’s efficiency mainly comes from its efficient model design, and that it requires no classical post-processing during training and can be entire trained and executed on GPU. Further, TELGEN only processes SD pairs with actual traffic demands, significantly reducing its memory footprint and improving training and prediction efficiency. TEAL’s fairly complex model pipeline (combining GNN with MARL) and the need for the expensive ADMM post-processing contributed mainly to its inefficiency, especially since ADMM can only be run on CPU. HARP has a complex model architecture with set transformers and specially designed recurrent adjustment units (RAUs), which contributed to its slow training and prediction. Both TEAL and HARP also process full traffic matrices regardless of whether a pair of nodes have demands or not, further increasing memory footprint and slowing down training and prediction.
@@
While the above results show TELGEN’s superior performance and efficiency compared to state-of-the-art algorithms when training and testing on the same topology, the next sets of experiments fully unveil TELGEN’s generalization power, for which no existing learning-based TE solution can compare. As a remark, while HARP also has some limited generalizability, it is hard to fairly compare TELGEN’s and HARP’s generalization power for two reasons: (1) their different objectives make their solutions not easily comparable, as shown in Table II; (2) HARP only adapts to up to $20\%$ differences from the training topology as in [3], which is not comparable to the goal of TELGEN to generalize to totally unseen, potentially much larger topologies during testing compared to training. Therefore, we next compare TELGEN’s performance to the most generalizable baseline—the SciPy IPM solver itself.
@@ V-D2 Generalization on Real-world Networks
The results are shown in Table IV.
@@
For the two ASN test sets, note that the training set contained only two much smaller subnetworks of the entire network, with around $40\%$ of the links as in $\text{ASN}^{\sf test}_{553}$ and $7\%$ of the links as in $\text{ASN}^{\sf test}_{1739}$. Also note that the training set had no overlap in nodes or links with the $\text{ASN}^{\sf test}_{553}$ due to how we constructed the training and test sets. Nevertheless, we observe that TELGEN achieved almost negligible OGap ($<2.5\%$) and CGap ($<0.65\%$) in both test sets compared to the optimal solution output by IPM. After scaling, the OnoCGap was $<2.1\%$. The OnoCGap was reduced compared to OGap since TELGEN tended to “slightly over-provision”—allocating more than link capacities which led to (infeasibly) higher objective than the optimal solution—which was corrected after the solution was scaled down by capacity violation. This was precisely the behavior we wanted TELGEN to learn: to maximize the objective as the IPM algorithm did even when the “maximize” symbol itself was not encoded in the GNN inputs at all.
TELGEN achieved $16.5\%$-$71.2\%$ reduction in prediction time over SciPy. The advantage is expected to become larger with larger TE instances (e.g. larger networks and more SD pairs). In principle, TELGEN involves linear complexity in the number of variables while the fastest LP solver involves almost cubic complexity. Moreover, TELGEN can utilize massive parallelism enabled by GPU, while traditional solvers can at best perform CPU parallelism, further limiting scalability.
TABLE V: Gaps (%) for all ER test datasets
Dataset OGap CGap OnoCGap
$\text{ER}^{\sf test}_{200,0.7}$ 1.36 0.11 1.25
$\text{ER}^{\sf test}_{200,0.8}$ 1.27 0.12 1.16
$\text{ER}^{\sf test}_{200,0.9}$ 0.97 0.09 0.73
$\text{ER}^{\sf test}_{500,0.7}$ 1.27 0.11 1.17
$\text{ER}^{\sf test}_{500,0.8}$ 1.29 0.11 1.20
$\text{ER}^{\sf test}_{500,0.9}$ 0.97 0.08 0.85
$\text{ER}^{\sf test}_{1000,0.7}$ 1.21 0.10 1.13
$\text{ER}^{\sf test}_{1000,0.8}$ 1.40 0.11 1.31
$\text{ER}^{\sf test}_{1000,0.9}$ 0.97 0.07 0.90
$\text{ER}^{\sf test}_{2000,0.7}$ 1.24 0.10 1.16
$\text{ER}^{\sf test}_{2000,0.8}$ 1.27 0.11 1.20
$\text{ER}^{\sf test}_{2000,0.9}$ 0.99 0.07 0.92
@@ V-D3 Generalization on Random Topologies
In random dataset experiments, we trained our TELGEN model on the 54 topology configurations for ER dataset and 7 topology configurations for WA dataset. The trained models had been tested on 12 test graphs with different number of nodes and probability for ER graphs while 4 test graphs with different number of nodes and $\beta$ for WA graphs. Table V and VI show test datasets with detailed node numbers and/or link generation parameters, as well as OGap, CGap and OnoCGap achieved for each test dataset. For both ER and WA datasets, TELGEN achieved negligible objective and constraint gaps, and the OnocGap was below $<2\%$ for all test sets. The increase in network sizes in the test datasets led to negligible impact on the gaps as well, demonstrating strong generalizability of TELGEN even when the test networks had $2$-20×20$\times$20 × more nodes than the largest training networks. The WA test set witnessed slightly larger gaps when the network became very larger with 6×6$\times$6 × more nodes than training, but the result was still superior.
For prediction time of generating a solution, Figs. 3 and 4 show that TELGEN again outperformed the traditional SciPy solver by over 3×3$\times$3 × for ER and WA test graphs. The more nodes or links the graph had, the more time TELGEN spent to make one prediction as observed in the figures. The advantage of TELGEN enlarged when the problem size—the number of SD pairs—increased as in Fig. 4(b). When the SD pair number went up to $500$, TELGEN sped up the traditional SciPy solver by over 200×200$\times$200 ×, validating the asymptotic advantage (linear vs. cubic) and demonstrating significant efficiency improvement for large-scale problems.
Figure 3: Prediction time for different ER test datasets based on number of nodes and edge probability
@@
Figure 4: Prediction time for different WA test datasets based on number of nodes and different number of SD pairs
V-D4 Generalization to Unseen Demand Distributions
Finally, we show TELGEN’s generalization power when given different ranges of demands than seen during training. We changed the demand range by $20\%$ over the original range of $[1000,5000]$, leading to two demand ranges: $[1200,6000]$ and $[800,4000]$. We tested the different demand ranges on the ER test datasets with a $0.9$ connection probability. Detailed information is shown in Table VII. TELGEN again achieved negligible gaps compared to the optimal solution given either larger or smaller demands than training, with the OnoCGap $<3\%$ for the former and $<0.8\%$ for the latter. The slightly higher gaps with higher demands were due to the link capacities being more saturated.
@@ V-D4 Generalization to Unseen Demand Distributions
$\text{ER}^{\sf test}_{200,0.9}$ [1200, 6000] 2.98 0.13 2.74
$\text{ER}^{\sf test}_{500,0.9}$ [1200, 6000] 3.09 0.13 2.89
$\text{ER}^{\sf test}_{1000,0.9}$ [1200, 6000] 2.66 0.13 2.45
$\text{ER}^{\sf test}_{2000,0.9}$ [1200, 6000] 2.94 0.13 2.73
$\text{ER}^{\sf test}_{200,0.9}$ [800, 4000] 0.75 0.09 0.71
$\text{ER}^{\sf test}_{500,0.9}$ [800, 4000] 0.79 0.09 0.75
$\text{ER}^{\sf test}_{1000,0.9}$ [800, 4000] 0.70 0.09 0.67
$\text{ER}^{\sf test}_{2000,0.9}$ [800, 4000] 0.75 0.09 0.71

### v2

@@ V-B2 TEAL [43]
We used the open-source implementation of TEAL with the original model architecture, hyperparameters and same training epochs as TELGEN to make fair comparison. As a remark, TEAL is designed in a way such that testing must be on the same topology and SD pairs as training.
@@ V-B3 HARP [2]
We used the open-source implementation of HARP with the original model architecture, hyperparameters and same training epochs as TELGEN to make fair comparison. One issue of HARP is that it has a large RAM footprint on large topologies, and even our latest hardware platform (as below) could not execute HARP on the largest topology—the full ASN. To make a fair comparison between TELGEN and the two baselines, we trained and tested all three algorithms above on B4 and a randomly sampled subnetwork of ASN in lieu of the full ASN topology.
@@ V-B4 SciPy [7]
V-B5 Gurobi [14]
We used the Gurobi Optimizer to obtain ground-truth optimal objectives for gap computation and also recorded its wall-clock solution time under default settings (LP barrier/simplex as chosen by the solver), without warm-starts.
Hardware platform: The algorithms were implemented and evaluated on a desktop server with Intel Xeon Gold 5317 CPU (3.0GHz, 48 cores), 256GB RAM and Nvidia A100 GPU (40GB). All experiments were run with $3$ random seeds and averages are shown below.
Runtime accounting
We report training time and per-instance inference latency instead of memory. TELGEN incurs a one-time training cost (about 4h), after which inference is millisecond-level. In contrast, classical solvers re-solve every instance from scratch. For a batch of $N$ instances, the amortized time is $T_{\text{TELGEN}}=T_{\text{train}}+N\cdot t_{\text{inf}}$ versus $T_{\text{solver}}=N\cdot t_{\text{solve}}$, favoring TELGEN for moderate $N$.
@@ V-C Evaluation Metrics
It is similar to the constraint loss but only considers final output of the algorithm. The objective gap without constraint violation (OnoCGap) is the objective gap when the solution is scaled by the maximum ratio of any constraint violation, thereby ensuring the final solution is feasible. Finally, the training time per epoch and the prediction time measure the time spent on training for each epoch on average, and the time for making each prediction on average during testing (or running IPM on one instance till optimality for SciPy), respectively.
V-D Evaluation Results
TABLE II: OnoCGap (%) for TEAL and TELGEN
@@ V-D Evaluation Results
V-D1 Comparing TELGEN, TEAL and HARP
We first compared the performance of TELGEN to TEAL and HARP, the state-of-the-art learning-based TE solutions. Since TEAL needs to be trained and tested on the same network topology while HARP cannot process topology as large as the whole ASN graph, we evaluated all three algorithms by training and testing them on datasets generated on B4 and a connected sub-network of ASN [^3] These are the smallest and largest topologies evaluated in [43] respectively, though we train and test on different SD pairs and demands than original paper. with $98$ nodes and $202$ egdes. Table II shows OnoCGap—the objective gap when all solutions are rounded to be feasible—of TELGEN, TEAL and HARP on the two datasets. To our surprise, TEAL performed poorly even on the small B4 topology, achieving as high as $26\%$ objective gap; on the sub-ASN topology, it had a $98\%$ objective gap. HARP performed almost as badly as TEAL as shown: it has a $27\%$ objective gap on B4 and $73\%$ on sub-ASN. As we investigate, the reason for TEAL’s inferior performance was due to lack of SD pairs for building the traffic matrix in the training dataset. The reasons for HARP’s performance were 1) the different objective of HARP leading to a different model architecture: HARP’s model architecture is designed to meet a specific goal which is maximum link utilization (MLU) in the paper while our TELGEN focuses on generalization across diverse topologies and network conditions, thus suits for different objectives; 2) In order to optimize for the MLU, HARP tended to allocate data flow through the few same nodes which caused bottleneck links and lead to the poor performance; 3) lack of SD pairs for building the traffic matrix in the training dataset which is similar as TEAL. Compared to TEAL and HARP, TELGEN achieved negligible ($3\%$) objective gap, and even performed better on the larger sub-ASN topology with $1.7\%$ gap after rounding.
Though HARP showed in [2] that it could handle topology variations, it only allowed $20\%$ of links capacity variation with ratio of less than 0.8, with about $5\%$ of the links with zero capacity (considering as inactive links) while TELGEN allows completely different topologies for training and testing. When considering topology changes for training and testing dataset, HARP only showed the effectiveness using AnonNet dataset with different sets of active nodes as clusters containing dozens of nodes but no larger graphs containing thousands or even more nodes like TELGEN. HARP also show the effectiveness in perturbations on a large topology, KDL with a 754 nodes, while TELGEN showed the generalization across different and much larger topology with the number of nodes up to $10^{6}$ as shown in the next section. This further showed the superior of our algorithm TELGEN. However, we are also aware that HARP is specially designed for an objective and the data we used might be not the best fit for HARP algorithm. We shall implement and adjust the HARP dataset so as to fit our TELGEN as a future work.
Both TEAL and HARP had comparable or faster training time per epoch as TELGEN on B4, but both became at least $3\times$ slower than TELGEN when facing the larger sub-ASN topology as shown in Table III. This was due to TEAL’s fairly complex model architecture which combines GNN with MARL, and even an expensive classical ADMM algorithm that must be run on CPU during end-to-end training. HARP’s model architecture could be even more complex that it starts with a GNN layer, after which comes a transformer as well as an MLP, then ended with recurrent adjustment unit (RAU) consisting layers of MLPs. TELGEN was faster because of its simple GNN architecture, and its entire training can be conducted on GPU. The reason that TEAL and HARP were slower than TELGEN is also because they both consider demands of all possible SD pairs among all nodes in the network, which means they take much larger size of inputs than TELGEN. On the contrary, TELGEN can be applied to an arbitrary number of SD pairs in the network which shows the flexibility of our method. For prediction time when making a prediction, TELGEN again significantly outperformed TEAL by over $3\times$ for B4 and sub-ASN, and HARP by over $18\times$ for B4 and $500\times$ for sub-ASN due to its simple, GPU-compatible architecture. The significant superior performance of TELGEN for training time and prediction time show that our model can be generalized to different SD pairs as well as topology sizes and be used with flexibility.
@@ V-D1 Comparing TELGEN, TEAL and HARP
While the above results show TELGEN’s superior performance and efficiency compared to state-of-the-art algorithm when training and testing on the same topology, the next sets of experiments fully unveil TELGEN’s generalization power, for which no existing learning-based TE solution can compare.
@@ V-D2 Generalization on Real-world Networks
We next show TELGEN’s generalization power on real-world networks, and compare its performance to the other generalizable baseline—the SciPy IPM solver itself. The results are shown in Table IV.
@@
The B4 results were the same as in Table II. For the two ASN test sets, note that the training set contained only two much smaller subnetworks of the entire network, with around $40\%$ of the links as in $\text{ASN}^{\sf test}_{553}$ and $7\%$ of the links as in $\text{ASN}^{\sf test}_{1739}$. Also note that the training set had no overlap in nodes or links with the $\text{ASN}^{\sf test}_{553}$ due to how we constructed the training and test sets. Nevertheless, we observe that TELGEN achieved almost negligible OGap ($2.5\%$) and CGap ($0.65\%$) in both test sets compared to the optimal solution output by IPM. After scaling, the OnoCGap was $2.1\%$. The OnoCGap was reduced compared to OGap since TELGEN tended to “over-provision”—allocating more than link capacities which led to (infeasibly) higher objective than the optimal solution—which was corrected after the solution was scaled down by capacity violation. This was exactly the behavior that we wanted TELGEN to learn: to maximize the objective as the IPM algorithm did even when the “maximize” symbol itself was not encoded in the GNN inputs at all.
For prediction time, TELGEN achieved $16.5\%$-$71.2\%$ reduction over SciPy. The advantage is expected to become larger with larger TE instances (e.g. larger networks and more SD pairs). In principle, TELGEN involves linear complexity in the number of variables while the fastest LP solver involves almost cubic complexity. Moreover, TELGEN can utilize the massive parallelism enabled by GPU, whereas traditional solvers can at best perform CPU parallelism, further limiting their scalability.
TABLE V: Gaps (%) for ER test datasets across varying densities
Dataset OGap (%) CGap (%) OnoCGap (%) Avg_inference_time (ms)
$\text{ER}^{\sf test}_{200,0.1}$ 3.88 6.14 0.30 $2.43\pm 0.87$
$\text{ER}^{\sf test}_{200,0.5}$ 3.31 5.17 0.48 $1.98\pm 1.26$
$\text{ER}^{\sf test}_{200,0.9}$ 3.18 5.48 0.58 $1.95\pm 1.27$
$\text{ER}^{\sf test}_{500,0.1}$ 3.71 5.25 0.41 $2.00\pm 1.26$
$\text{ER}^{\sf test}_{500,0.5}$ 2.90 5.23 0.44 $1.92\pm 1.29$
$\text{ER}^{\sf test}_{500,0.9}$ 3.48 5.53 0.49 $1.97\pm 1.28$
$\text{ER}^{\sf test}_{1000,0.1}$ 3.53 5.00 0.64 $1.90\pm 1.24$
$\text{ER}^{\sf test}_{1000,0.5}$ 3.34 5.21 0.48 $1.88\pm 1.88$
$\text{ER}^{\sf test}_{1000,0.9}$ 3.19 5.52 0.43 $1.92\pm 1.31$
$\text{ER}^{\sf test}_{2000,0.1}$ 3.82 5.02 0.71 $1.87\pm 1.22$
$\text{ER}^{\sf test}_{2000,0.5}$ 3.31 5.17 0.48 $1.97\pm 1.28$
$\text{ER}^{\sf test}_{2000,0.9}$ 3.18 5.48 0.58 $1.91\pm 1.22$
@@ V-D3 Generalization on Random Topologies
In random dataset experiments, we trained our TELGEN model on the 54 topology configurations for ER dataset and 7 topology configurations for WA dataset. The trained models had been tested on 12 test graphs with different number of nodes and probability for ER graphs while 4 test graphs with different number of nodes and $\beta$ for WA graphs. Table V and VI show the test dataset detailed node number and edge information as well as three kinds of gaps: OGap, CGap, OnoCGap for each test dataset. For both ER and WA datasets, TELGEN achieved negligible objective and constraint gaps, and the OnocGap was below $2\%$ for all test sets. The increase in network sizes in the test sets led to negligible impact on the gaps as well, demonstrating strong generalizability of TELGEN even when the test networks had $2$-$20\times$ more nodes than the largest training networks. The WA test set witnessed slightly larger gaps when the network became very larger with $6\times$ more nodes than training, but the result was still superior.
Fig. 3 reports the distribution of inference time across edge probability $p\in[0.1,0.9]$ and graph sizes $n\in\{200,500,1000,2000\}$. Among the remaining solvers, TELGEN (teal) consistently achieves a lower mean latency than both Scipy and Gurobi, while maintaining tight variability across all $p$ and $n$. Its runtime grows smoothly with $n$ and shows weak sensitivity to $p$, reflecting TELGEN’s fixed-depth, structure-aware forward pass; in contrast, classical LP/MIP pipelines require per-instance linear solves or search, resulting in higher averages runtime. Fig. 4 shows that TELGEN again outperforms traditional solvers, and the advantage enlarges as the problem size (the number of SD pairs) increases. When the SD pair count reaches $500$, TELGEN is over $200\times$ faster than the traditional Scipy solver, validating the asymptotic advantage (linear vs. cubic) and demonstrating significant efficiency for large-scale problems.
The advantage of TELGEN enlarged when the problem size—the number of SD pairs—increased as in Fig. 4(b). When the SD pair number went up to $500$, TELGEN sped up the traditional SciPy solver by over $200\times$, validating the asymptotic advantage (linear vs. cubic) and demonstrating significant efficiency improvement for large-scale problems.
Figure 3: Inference time (ms) across edge probability $p$ for ER graphs, faceted by size $n\in\{200,500,1000,2000\}$.
TELGEN completes a one-time training phase (around four hours) and subsequently performs inference within milliseconds. By contrast, Scipy re-solves each instance without reuse of prior computations, resulting in significantly higher average computation delay and frequent outliers across edge probabilities and graph scales, as illustrated in Fig. 3.
@@
Figure 4: Prediction time for different WA test datasets based on number of nodes and different number of SD pairs.
@@
$\text{ER}^{\sf test}_{200,0.9}$ [1200, 6000] 3.01 0.15 2.85
$\text{ER}^{\sf test}_{500,0.9}$ [1200, 6000] 2.99 0.15 2.86
$\text{ER}^{\sf test}_{1000,0.9}$ [1200, 6000] 3.00 0.15 2.84
$\text{ER}^{\sf test}_{2000,0.9}$ [1200, 6000] 3.02 0.15 2.86
$\text{ER}^{\sf test}_{200,0.9}$ [800, 4000] 3.02 0.15 2.86
$\text{ER}^{\sf test}_{500,0.9}$ [800, 4000] 2.99 0.15 2.86
$\text{ER}^{\sf test}_{1000,0.9}$ [800, 4000] 3.01 0.15 2.85
$\text{ER}^{\sf test}_{2000,0.9}$ [800, 4000] 3.00 0.15 2.84
V-D4 Generalization to Unseen Demand Distributions
We evaluate TELGEN on Erdős–Rényi (ER) test graphs of different scales and sparsities, using node counts $n\!\in\!\{200,500,1000,2000\}$ and connection probabilities $p$ from $0.1$ to $0.9$. As reported in Table V, TELGEN maintains consistently small errors across all settings: OGap ranges 3.18–3.88%, CGap ranges 5–6.14%, and the stricter OnoCGap remains within 0.30–0.71%. The per-instance prediction time is low and stable, with means in 1.87–2.43 ms and standard deviations 0.87–1.88 ms, showing weak dependence on $n$ or $p$. As shown in Table VII, the gaps remain nearly identical across all unseen demand distributions, demonstrating TELGEN’s strong robustness to varying demand ranges. The OGap and OnoCGap values fluctuate within a narrow margin of ±0.02–0.03 %, while CGap stays constant at 0.15 %. Notably, the performance is invariant to both network size (200–2000 nodes) and demand scale ([800, 4000] vs. [1200, 6000]), confirming that TELGEN generalizes well to unseen traffic intensities without degradation in optimality. These results indicate that TELGEN generalizes well across graph scale and density while preserving near-optimality and fast inference.